OpenStudy (anonymous):

Every pay period, samuel budgets 1/2 of his paycheck for expenses, 1/6 for recreation , and the rest for savings. his paycheck his $894.30. How much money does samuel put aside for savings each pay period ? *help and i'll give medal*

OpenStudy (anonymous):

this one is a bit hard xD

OpenStudy (anonymous):

thats wut she said o-0

OpenStudy (anonymous):

xD haha lol

OpenStudy (anonymous):

so do you know it or nah

OpenStudy (anonymous):

neh

OpenStudy (anonymous):

im kinda lost xD

OpenStudy (anonymous):

lol okay

OpenStudy (blackops2luvr):

:3

OpenStudy (anonymous):

\[\left(1-\left(\frac{1}{2}+\frac{1}{6}\right)\right) 894.3=298.1 \]

OpenStudy (anonymous):

i should have said what i thought it was .-.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

cuz i would have gotten it right o-0

OpenStudy (blackops2luvr):

\(\bbox[5pt,#E50000]{\!} \bbox[5pt,#E53600]{\!} \bbox[5pt,#E56000]{\!} \bbox[5pt,#E58D00]{\!} \bbox[5pt,#E5B600]{\!} \bbox[5pt,#E5D600]{\!} \bbox[5pt,#D0E500]{\!} \bbox[5pt,#B0E500]{\!} \bbox[5pt,#8FE500]{\!} \bbox[5pt,#6EE500]{\!} \bbox[5pt,#41E500]{\!} \bbox[5pt,#01E500]{\!} \bbox[5pt,#00E564]{\!} \bbox[5pt,#00E5AD]{\!} \bbox[5pt,#00E5C9]{\!} \bbox[5pt,#00E0E5]{\!} \bbox[5pt,#00CFE5]{\!} \bbox[5pt,#00B9E5]{\!} \bbox[5pt,#009FE5]{\!} \bbox[5pt,#0087E5]{\!} \bbox[5pt,#0074E5]{\!} \bbox[5pt,#005AE5]{\!} \bbox[5pt,#003DE5]{\!} \bbox[5pt,#002CE5]{\!} \bbox[5pt,#0600E5]{\!} \bbox[5pt,#1E00E5]{\!} \bbox[5pt,#2C00E5]{\!} \bbox[5pt,#3B00E5]{\!} \bbox[5pt,#5000E5]{\!} \bbox[5pt,#6500E5]{\!} \bbox[5pt,#8200E5]{\!} \bbox[5pt,#A000E5]{\!} \bbox[5pt,#BA00E5]{\!} \bbox[5pt,#CE00E5]{\!} \bbox[5pt,#E400E5]{\!} \bbox[5pt,#E500C7]{\!} \bbox[5pt,#E500AC]{\!} \bbox[5pt,#E50093]{\!} \bbox[5pt,#E50077]{\!} \bbox[5pt,#E5005C]{\!} \bbox[5pt,#E50039]{\!} \bbox[5pt,#E5001C]{\!}\)

OpenStudy (blackops2luvr):

MAH BABY I STILL HAVE YOU IN MY DOCS :D

OpenStudy (blackops2luvr):

i made tht last year XD

OpenStudy (anonymous):

Haha alright i gotta go on to a different thing gotta get this done

OpenStudy (blackops2luvr):

'.'

OpenStudy (blackops2luvr):

Fishhyyss

OpenStudy (blackops2luvr):

:D

OpenStudy (blackops2luvr):



OpenStudy (blackops2luvr):

^ worst thing ever ^

OpenStudy (anonymous):